# arXiv:2603.01962v1[quant-ph] 02 Mar 2026

### v1

@@ I.2 Dynamic Bayesian network approach (DBN)
$p(e_{\mathrm{e}}^{j}|\lambda_{1}^{\alpha})$ $=\text{tr}\left[\Pi_{\mathrm{e}}^{j}\,\frac{P_{1}^{\alpha}\tilde{\rho}_{1}P_{1}^{\alpha}}{\text{tr}(P_{1}^{\alpha}\tilde{\rho}_{1})}\right],$ (S24)
$p(e_{\mathrm{k}}^{l}|\lambda_{2}^{\beta})$ $=\text{tr}\left[\Pi_{\mathrm{k}}^{l}\,\frac{P_{2}^{\beta}\tilde{\rho}_{2}P_{2}^{\beta}}{\text{tr}(P_{2}^{\beta}\tilde{\rho}_{2})}\right],$
$p(e_{\mathrm{k}}^{m}|\lambda_{3}^{\gamma})$ $=\text{tr}\left[\Pi_{\mathrm{k}}^{m}\,\frac{P_{3}^{\gamma}\tilde{\rho}_{3}P_{3}^{\gamma}}{\text{tr}(P_{3}^{\gamma}\tilde{\rho}_{3})}\right],$
$p(e_{\mathrm{e}}^{n}|\lambda_{4}^{\delta})$ $=\text{tr}\left[\Pi_{\mathrm{e}}^{n}\,\frac{P_{4}^{\delta}\tilde{\rho}_{4}P_{4}^{\delta}}{\text{tr}(P_{4}^{\delta}\tilde{\rho}_{4})}\right],$
$p(e_{\mathrm{e}}^{r}|\lambda_{1}^{\epsilon})$ $=\text{tr}\left[\Pi_{\mathrm{e}}^{r}\,\frac{P_{1}^{\epsilon}\tilde{\rho}_{1}P_{1}^{\epsilon}}{\text{tr}(P_{1}^{\epsilon}\tilde{\rho}_{1})}\right].$
@@
$p_{w}^{\mathrm{DBN}}(e_{\mathrm{e}}^{j},e_{\mathrm{k}}^{l},e_{\mathrm{k}}^{m},e_{\mathrm{e}}^{n})=$ $\sum_{\alpha,\beta,\gamma,\delta}\text{tr}(P_{1}^{\alpha}\tilde{\rho}_{1})\,\text{tr}\left\{P_{2}^{\beta}\,\mathcal{U}_{\mathrm{k}}\left[\frac{P_{1}^{\alpha}\tilde{\rho}_{1}P_{1}^{\alpha}}{\text{tr}(P_{1}^{\alpha}\tilde{\rho}_{1})}\right]\right\}\,\text{tr}\left\{P_{3}^{\gamma}\,\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\left[\frac{P_{2}^{\beta}\tilde{\rho}_{2}P_{2}^{\beta}}{\text{tr}(P_{2}^{\beta}\tilde{\rho}_{2})}\right]\right\}\,\text{tr}\left\{P_{4}^{\delta}\,\mathcal{U}_{\mathrm{e}}\left[\frac{P_{3}^{\gamma}\tilde{\rho}_{3}P_{3}^{\gamma}}{\text{tr}(P_{3}^{\gamma}\tilde{\rho}_{3})}\right]\right\},$ (S26)
$\times\text{tr}\left[\Pi_{\mathrm{e}}^{j}\,\frac{P_{1}^{\alpha}\tilde{\rho}_{1}P_{1}^{\alpha}}{\text{tr}(P_{1}^{\alpha}\tilde{\rho}_{1})}\right]\,\text{tr}\left[\Pi_{\mathrm{k}}^{l}\,\frac{P_{2}^{\beta}\tilde{\rho}_{2}P_{2}^{\beta}}{\text{tr}(P_{2}^{\beta}\tilde{\rho}_{2})}\right]\,\text{tr}\left[\Pi_{\mathrm{k}}^{m}\,\frac{P_{3}^{\gamma}\tilde{\rho}_{3}P_{3}^{\gamma}}{\text{tr}(P_{3}^{\gamma}\tilde{\rho}_{3})}\right]\,\text{tr}\left[\Pi_{\mathrm{e}}^{n}\,\frac{P_{4}^{\delta}\tilde{\rho}_{4}P_{4}^{\delta}}{\text{tr}(P_{4}^{\delta}\tilde{\rho}_{4})}\right],$
$p_{q_{\mathrm{h}}}^{\mathrm{DBN}}(e_{\mathrm{k}}^{l},e_{\mathrm{k}}^{m})=$ $\sum_{\beta,\gamma}\text{tr}\left[P_{3}^{\gamma}\,\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\left(P_{2}^{\beta}\tilde{\rho}_{2}P_{2}^{\beta}\right)\right]\,\text{tr}\left[\Pi_{\mathrm{k}}^{l}\,\frac{P_{2}^{\beta}\tilde{\rho}_{2}P_{2}^{\beta}}{\text{tr}(P_{2}^{\beta}\tilde{\rho}_{2})}\right]\,\text{tr}\left[\Pi_{\mathrm{k}}^{m}\,\frac{P_{3}^{\gamma}\tilde{\rho}_{3}P_{3}^{\gamma}}{\text{tr}(P_{3}^{\gamma}\tilde{\rho}_{3})}\right],$
$p_{q_{\mathrm{c}}}^{\mathrm{DBN}}(e_{\mathrm{e}}^{n},e_{\mathrm{e}}^{r})=$ $\sum_{\delta,\epsilon}\text{tr}\left[P_{1}^{\epsilon}\,\mathcal{T}_{t_{\mathrm{c}}}^{\mathrm{c}}\left(P_{4}^{\delta}\tilde{\rho}_{4}P_{4}^{\delta}\right)\right]\,\text{tr}\left[\Pi_{\mathrm{e}}^{n}\,\frac{P_{4}^{\delta}\tilde{\rho}_{4}P_{4}^{\delta}}{\text{tr}(P_{4}^{\delta}\tilde{\rho}_{4})}\right]\,\text{tr}\left[\Pi_{\mathrm{e}}^{r}\,\frac{P_{1}^{\epsilon}\tilde{\rho}_{1}P_{1}^{\epsilon}}{\text{tr}(P_{1}^{\epsilon}\tilde{\rho}_{1})}\right].$
@@
$\sigma_{w}^{2}$ $=\sigma_{q_{\mathrm{h}}}^{2}+\sigma_{q_{\mathrm{c}}}^{2}-2\langle q_{\mathrm{h}}\rangle\langle q_{\mathrm{c}}\rangle+2\sum_{\delta,\epsilon}\text{tr}\left[P_{1}^{\epsilon}\,\mathcal{T}_{t_{\mathrm{c}}}^{\mathrm{c}}\left(P_{4}^{\delta}\tilde{\rho}_{4}P_{4}^{\delta}\right)\right]\,\text{tr}\left[H_{\mathrm{e}}\,\frac{P_{4}^{\delta}\tilde{\rho}_{4}P_{4}^{\delta}}{\text{tr}(P_{4}^{\delta}\tilde{\rho}_{4})}\right]\,\text{tr}\left[H_{\mathrm{e}}\,\frac{P_{1}^{\epsilon}\tilde{\rho}_{1}P_{1}^{\epsilon}}{\text{tr}(P_{1}^{\epsilon}\tilde{\rho}_{1})}\right]$ (S28)
@@
$-2\sum_{\alpha,\beta,\gamma,\delta}\text{tr}(P_{1}^{\alpha}\tilde{\rho}_{1})\,\text{tr}\left\{P_{2}^{\beta}\,\mathcal{U}_{\mathrm{k}}\left[\frac{P_{1}^{\alpha}\tilde{\rho}_{1}P_{1}^{\alpha}}{\text{tr}(P_{1}^{\alpha}\tilde{\rho}_{1})}\right]\right\}\,\text{tr}\left\{P_{3}^{\gamma}\,\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\left[\frac{P_{2}^{\beta}\tilde{\rho}_{2}P_{2}^{\beta}}{\text{tr}(P_{2}^{\beta}\tilde{\rho}_{2})}\right]\right\}\,\text{tr}\left\{P_{4}^{\delta}\,\mathcal{U}_{\mathrm{e}}\left[\frac{P_{3}^{\gamma}\tilde{\rho}_{3}P_{3}^{\gamma}}{\text{tr}(P_{3}^{\gamma}\tilde{\rho}_{3})}\right]\right\}$
@@
$\times\text{tr}\left[H_{\mathrm{e}}\,\frac{P_{1}^{\alpha}\tilde{\rho}_{1}P_{1}^{\alpha}}{\text{tr}(P_{1}^{\alpha}\tilde{\rho}_{1})}\right]\,\text{tr}\left[H_{\mathrm{e}}\,\frac{P_{4}^{\delta}\tilde{\rho}_{4}P_{4}^{\delta}}{\text{tr}(P_{4}^{\delta}\tilde{\rho}_{4})}\right]$
@@
$+2\sum_{\beta,\gamma,\delta}\text{tr}\left[P_{3}^{\gamma}\,\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\left(P_{2}^{\beta}\tilde{\rho}_{2}P_{2}^{\beta}\right)\right]\,\text{tr}\left\{P_{4}^{\delta}\,\mathcal{U}_{\mathrm{e}}\left[\frac{P_{3}^{\gamma}\tilde{\rho}_{3}P_{3}^{\gamma}}{\text{tr}(P_{3}^{\gamma}\tilde{\rho}_{3})}\right]\right\}\,\text{tr}\left[H_{\mathrm{k}}\,\frac{P_{2}^{\beta}\tilde{\rho}_{2}P_{2}^{\beta}}{\text{tr}(P_{2}^{\beta}\tilde{\rho}_{2})}\right]\,\text{tr}\left[H_{\mathrm{e}}\,\frac{P_{4}^{\delta}\tilde{\rho}_{4}P_{4}^{\delta}}{\text{tr}(P_{4}^{\delta}\tilde{\rho}_{4})}\right]$
@@
$+2\sum_{\alpha,\beta,\gamma}\text{tr}(P_{1}^{\alpha}\tilde{\rho}_{1})\,\text{tr}\left\{P_{2}^{\beta}\,\mathcal{U}_{\mathrm{k}}\left[\frac{P_{1}^{\alpha}\tilde{\rho}_{1}P_{1}^{\alpha}}{\text{tr}(P_{1}^{\alpha}\tilde{\rho}_{1})}\right]\right\}\,\text{tr}\left\{P_{3}^{\gamma}\,\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\left[\frac{P_{2}^{\beta}\tilde{\rho}_{2}P_{2}^{\beta}}{\text{tr}(P_{2}^{\beta}\tilde{\rho}_{2})}\right]\right\}\,\text{tr}\left[H_{\mathrm{e}}\,\frac{P_{1}^{\alpha}\tilde{\rho}_{1}P_{1}^{\alpha}}{\text{tr}(P_{1}^{\alpha}\tilde{\rho}_{1})}\right]\,\text{tr}\left[H_{\mathrm{k}}\,\frac{P_{3}^{\gamma}\tilde{\rho}_{3}P_{3}^{\gamma}}{\text{tr}(P_{3}^{\gamma}\tilde{\rho}_{3})}\right]$
@@
$-2\sum_{\alpha,\beta}\text{tr}(P_{1}^{\alpha}\tilde{\rho}_{1})\,\text{tr}\left\{P_{2}^{\beta}\,\mathcal{U}_{\mathrm{k}}\left[\frac{P_{1}^{\alpha}\tilde{\rho}_{1}P_{1}^{\alpha}}{\text{tr}(P_{1}^{\alpha}\tilde{\rho}_{1})}\right]\right\}\,\text{tr}\left[H_{\mathrm{e}}\,\frac{P_{1}^{\alpha}\tilde{\rho}_{1}P_{1}^{\alpha}}{\text{tr}(P_{1}^{\alpha}\tilde{\rho}_{1})}\right]\,\text{tr}\left[H_{\mathrm{k}}\,\frac{P_{2}^{\beta}\tilde{\rho}_{2}P_{2}^{\beta}}{\text{tr}(P_{2}^{\beta}\tilde{\rho}_{2})}\right]$
@@
$-2\sum_{\gamma,\delta}\text{tr}\left[P_{4}^{\delta}\,\mathcal{U}_{\mathrm{e}}\left(P_{3}^{\gamma}\tilde{\rho}_{3}P_{3}^{\gamma}\right)\right]\,\text{tr}\left[H_{\mathrm{k}}\,\frac{P_{3}^{\gamma}\tilde{\rho}_{3}P_{3}^{\gamma}}{\text{tr}(P_{3}^{\gamma}\tilde{\rho}_{3})}\right]\,\text{tr}\left[H_{\mathrm{e}}\,\frac{P_{4}^{\delta}\tilde{\rho}_{4}P_{4}^{\delta}}{\text{tr}(P_{4}^{\delta}\tilde{\rho}_{4})}\right].$
@@
$+2\,\text{tr}(H_{\mathrm{k}}\tilde{\rho}_{3})\text{tr}(H_{\mathrm{k}}\tilde{\rho}_{2})-2\sum_{\beta,\gamma}\text{tr}\left[P_{3}^{\gamma}\,\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\left(P_{2}^{\beta}\tilde{\rho}_{2}P_{2}^{\beta}\right)\right]\,\text{tr}\left[H_{\mathrm{k}}\,\frac{P_{2}^{\beta}\tilde{\rho}_{2}P_{2}^{\beta}}{\text{tr}(P_{2}^{\beta}\tilde{\rho}_{2})}\right]\,\text{tr}\left[H_{\mathrm{k}}\,\frac{P_{3}^{\gamma}\tilde{\rho}_{3}P_{3}^{\gamma}}{\text{tr}(P_{3}^{\gamma}\tilde{\rho}_{3})}\right].$
@@
$+2\,\text{tr}(H_{\mathrm{e}}\tilde{\rho}_{1})\text{tr}(H_{\mathrm{e}}\tilde{\rho}_{4})-2\sum_{\delta,\epsilon}\text{tr}\left[P_{1}^{\epsilon}\,\mathcal{T}_{t_{\mathrm{c}}}^{\mathrm{c}}\left(P_{4}^{\delta}\tilde{\rho}_{4}P_{4}^{\delta}\right)\right]\,\text{tr}\left[H_{\mathrm{e}}\,\frac{P_{4}^{\delta}\tilde{\rho}_{4}P_{4}^{\delta}}{\text{tr}(P_{4}^{\delta}\tilde{\rho}_{4})}\right]\,\text{tr}\left[H_{\mathrm{e}}\,\frac{P_{1}^{\epsilon}\tilde{\rho}_{1}P_{1}^{\epsilon}}{\text{tr}(P_{1}^{\epsilon}\tilde{\rho}_{1})}\right].$

### v2

@@ I.2 Dynamic Bayesian network approach (DBN)
$p(e_{\mathrm{e}}^{j}|\lambda_{1}^{\alpha})$ $=\text{tr}\left[\Pi_{\mathrm{e}}^{j}\,\frac{\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1}\mathrm{P}_{1}^{\alpha}}{\text{tr}(\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1})}\right],$ (S24)
$p(e_{\mathrm{k}}^{l}|\lambda_{2}^{\beta})$ $=\text{tr}\left[\Pi_{\mathrm{k}}^{l}\,\frac{\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2}\mathrm{P}_{2}^{\beta}}{\text{tr}(\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2})}\right],$
$p(e_{\mathrm{k}}^{m}|\lambda_{3}^{\gamma})$ $=\text{tr}\left[\Pi_{\mathrm{k}}^{m}\,\frac{\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3}\mathrm{P}_{3}^{\gamma}}{\text{tr}(\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3})}\right],$
$p(e_{\mathrm{e}}^{n}|\lambda_{4}^{\delta})$ $=\text{tr}\left[\Pi_{\mathrm{e}}^{n}\,\frac{\mathrm{P}_{4}^{\delta}\tilde{\rho}_{4}\mathrm{P}_{4}^{\delta}}{\text{tr}(\mathrm{P}_{4}^{\delta}\tilde{\rho}_{4})}\right],$
$p(e_{\mathrm{e}}^{r}|\lambda_{1}^{\epsilon})$ $=\text{tr}\left[\Pi_{\mathrm{e}}^{r}\,\frac{\mathrm{P}_{1}^{\epsilon}\tilde{\rho}_{1}\mathrm{P}_{1}^{\epsilon}}{\text{tr}(\mathrm{P}_{1}^{\epsilon}\tilde{\rho}_{1})}\right].$
@@
$p_{w}^{\mathrm{DBN}}(e_{\mathrm{e}}^{j},e_{\mathrm{k}}^{l},e_{\mathrm{k}}^{m},e_{\mathrm{e}}^{n})=$ $\sum_{\alpha,\beta,\gamma,\delta}\text{tr}(\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1})\,\text{tr}\left\{\mathrm{P}_{2}^{\beta}\,\mathcal{U}_{\mathrm{k}}\left[\frac{\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1}\mathrm{P}_{1}^{\alpha}}{\text{tr}(\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1})}\right]\right\}\,\text{tr}\left\{\mathrm{P}_{3}^{\gamma}\,\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\left[\frac{\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2}\mathrm{P}_{2}^{\beta}}{\text{tr}(\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2})}\right]\right\}\,\text{tr}\left\{\mathrm{P}_{4}^{\delta}\,\mathcal{U}_{\mathrm{e}}\left[\frac{\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3}\mathrm{P}_{3}^{\gamma}}{\text{tr}(\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3})}\right]\right\},$ (S26)
$\times\text{tr}\left[\Pi_{\mathrm{e}}^{j}\,\frac{\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1}\mathrm{P}_{1}^{\alpha}}{\text{tr}(\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1})}\right]\,\text{tr}\left[\Pi_{\mathrm{k}}^{l}\,\frac{\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2}\mathrm{P}_{2}^{\beta}}{\text{tr}(\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2})}\right]\,\text{tr}\left[\Pi_{\mathrm{k}}^{m}\,\frac{\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3}\mathrm{P}_{3}^{\gamma}}{\text{tr}(\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3})}\right]\,\text{tr}\left[\Pi_{\mathrm{e}}^{n}\,\frac{\mathrm{P}_{4}^{\delta}\tilde{\rho}_{4}\mathrm{P}_{4}^{\delta}}{\text{tr}(\mathrm{P}_{4}^{\delta}\tilde{\rho}_{4})}\right],$
$p_{q_{\mathrm{h}}}^{\mathrm{DBN}}(e_{\mathrm{k}}^{l},e_{\mathrm{k}}^{m})=$ $\sum_{\beta,\gamma}\text{tr}\left[\mathrm{P}_{3}^{\gamma}\,\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\left(\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2}\mathrm{P}_{2}^{\beta}\right)\right]\,\text{tr}\left[\Pi_{\mathrm{k}}^{l}\,\frac{\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2}\mathrm{P}_{2}^{\beta}}{\text{tr}(\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2})}\right]\,\text{tr}\left[\Pi_{\mathrm{k}}^{m}\,\frac{\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3}\mathrm{P}_{3}^{\gamma}}{\text{tr}(\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3})}\right],$
$p_{q_{\mathrm{c}}}^{\mathrm{DBN}}(e_{\mathrm{e}}^{n},e_{\mathrm{e}}^{r})=$ $\sum_{\delta,\epsilon}\text{tr}\left[\mathrm{P}_{1}^{\epsilon}\,\mathcal{T}_{t_{\mathrm{c}}}^{\mathrm{c}}\left(\mathrm{P}_{4}^{\delta}\tilde{\rho}_{4}\mathrm{P}_{4}^{\delta}\right)\right]\,\text{tr}\left[\Pi_{\mathrm{e}}^{n}\,\frac{\mathrm{P}_{4}^{\delta}\tilde{\rho}_{4}\mathrm{P}_{4}^{\delta}}{\text{tr}(\mathrm{P}_{4}^{\delta}\tilde{\rho}_{4})}\right]\,\text{tr}\left[\Pi_{\mathrm{e}}^{r}\,\frac{\mathrm{P}_{1}^{\epsilon}\tilde{\rho}_{1}\mathrm{P}_{1}^{\epsilon}}{\text{tr}(\mathrm{P}_{1}^{\epsilon}\tilde{\rho}_{1})}\right].$
@@
$\sigma_{w}^{2}$ $=\sigma_{q_{\mathrm{h}}}^{2}+\sigma_{q_{\mathrm{c}}}^{2}-2\langle q_{\mathrm{h}}\rangle\langle q_{\mathrm{c}}\rangle+2\sum_{\delta,\epsilon}\text{tr}\left[\mathrm{P}_{1}^{\epsilon}\,\mathcal{T}_{t_{\mathrm{c}}}^{\mathrm{c}}\left(\mathrm{P}_{4}^{\delta}\tilde{\rho}_{4}\mathrm{P}_{4}^{\delta}\right)\right]\,\text{tr}\left[H_{\mathrm{e}}\,\frac{\mathrm{P}_{4}^{\delta}\tilde{\rho}_{4}\mathrm{P}_{4}^{\delta}}{\text{tr}(\mathrm{P}_{4}^{\delta}\tilde{\rho}_{4})}\right]\,\text{tr}\left[H_{\mathrm{e}}\,\frac{\mathrm{P}_{1}^{\epsilon}\tilde{\rho}_{1}\mathrm{P}_{1}^{\epsilon}}{\text{tr}(\mathrm{P}_{1}^{\epsilon}\tilde{\rho}_{1})}\right]$ (S28)
@@
$-2\sum_{\alpha,\beta,\gamma,\delta}\text{tr}(\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1})\,\text{tr}\left\{\mathrm{P}_{2}^{\beta}\,\mathcal{U}_{\mathrm{k}}\left[\frac{\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1}\mathrm{P}_{1}^{\alpha}}{\text{tr}(\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1})}\right]\right\}\,\text{tr}\left\{\mathrm{P}_{3}^{\gamma}\,\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\left[\frac{\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2}\mathrm{P}_{2}^{\beta}}{\text{tr}(\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2})}\right]\right\}\,\text{tr}\left\{\mathrm{P}_{4}^{\delta}\,\mathcal{U}_{\mathrm{e}}\left[\frac{\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3}\mathrm{P}_{3}^{\gamma}}{\text{tr}(\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3})}\right]\right\}$
@@
$\times\text{tr}\left[H_{\mathrm{e}}\,\frac{\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1}\mathrm{P}_{1}^{\alpha}}{\text{tr}(\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1})}\right]\,\text{tr}\left[H_{\mathrm{e}}\,\frac{\mathrm{P}_{4}^{\delta}\tilde{\rho}_{4}\mathrm{P}_{4}^{\delta}}{\text{tr}(\mathrm{P}_{4}^{\delta}\tilde{\rho}_{4})}\right]$
@@
$+2\sum_{\beta,\gamma,\delta}\text{tr}\left[\mathrm{P}_{3}^{\gamma}\,\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\left(\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2}\mathrm{P}_{2}^{\beta}\right)\right]\,\text{tr}\left\{\mathrm{P}_{4}^{\delta}\,\mathcal{U}_{\mathrm{e}}\left[\frac{\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3}\mathrm{P}_{3}^{\gamma}}{\text{tr}(\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3})}\right]\right\}\,\text{tr}\left[H_{\mathrm{k}}\,\frac{\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2}\mathrm{P}_{2}^{\beta}}{\text{tr}(\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2})}\right]\,\text{tr}\left[H_{\mathrm{e}}\,\frac{\mathrm{P}_{4}^{\delta}\tilde{\rho}_{4}\mathrm{P}_{4}^{\delta}}{\text{tr}(\mathrm{P}_{4}^{\delta}\tilde{\rho}_{4})}\right]$
@@
$+2\sum_{\alpha,\beta,\gamma}\text{tr}(\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1})\,\text{tr}\left\{\mathrm{P}_{2}^{\beta}\,\mathcal{U}_{\mathrm{k}}\left[\frac{\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1}\mathrm{P}_{1}^{\alpha}}{\text{tr}(\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1})}\right]\right\}\,\text{tr}\left\{\mathrm{P}_{3}^{\gamma}\,\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\left[\frac{\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2}\mathrm{P}_{2}^{\beta}}{\text{tr}(\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2})}\right]\right\}\,\text{tr}\left[H_{\mathrm{e}}\,\frac{\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1}\mathrm{P}_{1}^{\alpha}}{\text{tr}(\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1})}\right]\,\text{tr}\left[H_{\mathrm{k}}\,\frac{\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3}\mathrm{P}_{3}^{\gamma}}{\text{tr}(\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3})}\right]$
@@
$-2\sum_{\alpha,\beta}\text{tr}(\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1})\,\text{tr}\left\{\mathrm{P}_{2}^{\beta}\,\mathcal{U}_{\mathrm{k}}\left[\frac{\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1}\mathrm{P}_{1}^{\alpha}}{\text{tr}(\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1})}\right]\right\}\,\text{tr}\left[H_{\mathrm{e}}\,\frac{\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1}\mathrm{P}_{1}^{\alpha}}{\text{tr}(\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1})}\right]\,\text{tr}\left[H_{\mathrm{k}}\,\frac{\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2}\mathrm{P}_{2}^{\beta}}{\text{tr}(\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2})}\right]$
@@
$-2\sum_{\gamma,\delta}\text{tr}\left[\mathrm{P}_{4}^{\delta}\,\mathcal{U}_{\mathrm{e}}\left(\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3}\mathrm{P}_{3}^{\gamma}\right)\right]\,\text{tr}\left[H_{\mathrm{k}}\,\frac{\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3}\mathrm{P}_{3}^{\gamma}}{\text{tr}(\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3})}\right]\,\text{tr}\left[H_{\mathrm{e}}\,\frac{\mathrm{P}_{4}^{\delta}\tilde{\rho}_{4}\mathrm{P}_{4}^{\delta}}{\text{tr}(\mathrm{P}_{4}^{\delta}\tilde{\rho}_{4})}\right].$
@@
$+2\,\text{tr}(H_{\mathrm{k}}\tilde{\rho}_{3})\text{tr}(H_{\mathrm{k}}\tilde{\rho}_{2})-2\sum_{\beta,\gamma}\text{tr}\left[\mathrm{P}_{3}^{\gamma}\,\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\left(\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2}\mathrm{P}_{2}^{\beta}\right)\right]\,\text{tr}\left[H_{\mathrm{k}}\,\frac{\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2}\mathrm{P}_{2}^{\beta}}{\text{tr}(\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2})}\right]\,\text{tr}\left[H_{\mathrm{k}}\,\frac{\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3}\mathrm{P}_{3}^{\gamma}}{\text{tr}(\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3})}\right].$
@@
$+2\,\text{tr}(H_{\mathrm{e}}\tilde{\rho}_{1})\text{tr}(H_{\mathrm{e}}\tilde{\rho}_{4})-2\sum_{\delta,\epsilon}\text{tr}\left[\mathrm{P}_{1}^{\epsilon}\,\mathcal{T}_{t_{\mathrm{c}}}^{\mathrm{c}}\left(\mathrm{P}_{4}^{\delta}\tilde{\rho}_{4}\mathrm{P}_{4}^{\delta}\right)\right]\,\text{tr}\left[H_{\mathrm{e}}\,\frac{\mathrm{P}_{4}^{\delta}\tilde{\rho}_{4}\mathrm{P}_{4}^{\delta}}{\text{tr}(\mathrm{P}_{4}^{\delta}\tilde{\rho}_{4})}\right]\,\text{tr}\left[H_{\mathrm{e}}\,\frac{\mathrm{P}_{1}^{\epsilon}\tilde{\rho}_{1}\mathrm{P}_{1}^{\epsilon}}{\text{tr}(\mathrm{P}_{1}^{\epsilon}\tilde{\rho}_{1})}\right].$